Bond A is a semi-annual coupon bond that has a face value of $1000, a 10% coupon rate, a five year maturity, and a yield to maturity of 7%. At the maturity date, how much payment should the bond investor expect from the bond? 


(a) $50 (b) $100 (c) $1035 (d) $1050
Answer #1

At maturity date the investor will receive coupon and par value
Amount received at maturity period = Par value + coupon Rate*Par Value/2 = 1000+ 1000*10%/2 = 1000 + 50 = 1050

Option d is correct option